if the roots of a quadratic equation are (-2+sqrt 6) and (-2-sqrt 6), what is the equation in ax^2+bx+c=0 form?
Found 3 solutions by solver91311, fcabanski, Onmyoji.S:
Answer by solver91311(24713)



If is a root of then is a factor of

So, multiply

John

Egw to Beta kai to Sigma
My calculator said it, I believe it, that settles it
The Out Campaign: Scarlet Letter of Atheism


Answer by fcabanski(1391)

If the roots of the equation were 3 and 5, the factors would be x-3 and x-5 and the equation would be the product of the factors = (x-3)(x-5) = x^2 -8x +15.


Subtract the roots from x to show the factors which are


x-(-2+sqrt(6)) and x - (-2-sqrt(6)) = x+2-sqrt(6) and x+2+sqrt(6). To find the equation multiply those factors together (multiply each term of the second factor by each term of the first factor).



Hope the solution helped. Sometimes you need more than a solution. Contact fcabanski@hotmail.com for online, private tutoring, or personalized problem solving (quick for groups of problems.)


Answer by Onmyoji.S(32)

if the roots of a quadratic equation are (-2+sqrt 6) and (-2-sqrt 6), what is the equation in ax^2+bx+c=0 form?
It would be:
(-2+sqrt 6)(-2-sqrt 6)
=4+2sqrt6-2sqrt6-(sqrt6)²
=4-(sqrt6)²